Write the equation of the line which passes through (2, –3) and is perpendicular to y = 4x + 7 in standard form. please show work.
Answer by ewatrrr(24785)

 
Hi,
Write the equation of the line which passes through (2, –3) and is perpendicular to y = 4x + 7 ( m = 4/1)
New Line: m = -1/4 (negative reciprocal of 4/1)
y = (-1/4)x + b
-3 = (-1/4)2 + b
-5/2 = b
y = (-1/4)x -7/2 or x + 4y = -14